SlimPay System Status
All Systems Operational
WEBSITE   Operational
PRODUCTION Operational
File API   Operational
Web API   Operational
Backoffice / Dashboard   Operational
Network   Operational
PREPROD Operational
File API   Operational
Web API   Operational
BackOffice / Dashboard   Operational
Network   Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Past Incidents
Nov 21, 2017

No incidents reported today.

Nov 20, 2017

No incidents reported.

Nov 19, 2017

No incidents reported.

Nov 18, 2017

No incidents reported.

Nov 17, 2017

No incidents reported.

Nov 16, 2017

No incidents reported.

Nov 15, 2017

No incidents reported.

Nov 14, 2017

No incidents reported.

Nov 13, 2017

No incidents reported.

Nov 12, 2017

No incidents reported.

Nov 11, 2017

No incidents reported.

Nov 10, 2017

No incidents reported.

Nov 9, 2017
Resolved - All services are back.
SMS services was impacted since 08h07 CET until 10h33 CET
Nov 9, 10:46 CET
Monitoring - The current major outage on OVH datacenter impact :
- our dev center (dev.slimpay.com) where our api documentation is exposed
- our sms service got some trouble (just got confirmation from our provider)
Nov 9, 10:42 CET
Nov 8, 2017

No incidents reported.

Nov 7, 2017
Resolved - Since now several days everything is running fine and all our metrics are green.
Nov 7, 22:34 CET
Monitoring - Analysis / Debrief :
During an DNS operation from our service provider an error was done. This DNS propagation was starting to impact SlimPay slowly around 07:10AM CET.
Infrastructure team started their investigation soon alert was received from SlimPay monitoring (from 7:13AM to 8:20AM). When we saw this issue was impacting all our client (based on the morning traffic) we contacted our service provider and got several exchange about it (between 8:30AM CET to 9:23AM CET). Following this discussion an incident was raised on SlimPay statuspage (9:23AM CET). Infrastructure team provided an temporary solution at 10:10AM CET in waiting a final solution from our service provider. At 11:52 our service provider found the root cause and applied an emergency fix.

Root cause: Our service provider did a DNS night operation which caused this service outage.

Next step: Until tomorrow we'll continue to use the temp solution tonight and tomorrow morning we'll see if the fix deployed by our service provider is well propagated before to rollback on it.
Nov 2, 17:57 CET
Update - Our infrastructure team has identified the root cause. It look like a DNS issue on our signature provider.
In waiting a resolution from our signature provider a temp fix has been set up by our infrastructure team. Service should be resumed now.
Every details (start time, end time, root cause, resolution) will be communicated in the next update
Nov 2, 10:10 CET
Investigating - We have detected some potential trouble on our signature provider. We are investigating on it and will provide update soon.
Nov 2, 09:23 CET